OpenStudy (anonymous):

Evaluate the velocity of a body and its distance moved, both after 2 minutes from it starts from rest, when its acceleration is given by a = 15 m/s^2

OpenStudy (anonymous):

Appreciate any help... Really stuck .. Thanks :)

OpenStudy (irishboy123):

this is constant acceleration, so you can use these equations of motion \[v = u + at \\ x = ut + \frac{1}{2} a t^2\] u, initial velocity = 0, starts "at rest"; v = velocity you are looking for x = distance travelled, you are also looking for t = time: careful with these units but, if you're actually supposed to be learning calculus, derive the equations from the givens.

OpenStudy (cwrw238):

acceleration = velocity / time so after 2 minutes (120 seconds) 15 = v / 120 so solve for v can you do that ?

OpenStudy (anonymous):

So for V will it be a(15) x t (120) = 1800 m/s ?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

So distance travelled is 108000m^2?

OpenStudy (cwrw238):

s = 1/2 a t^2 = 7.5 * 1800^2

OpenStudy (cwrw238):

since ut = 0 ( the object starts from rest)

OpenStudy (cwrw238):

these equations are for constant acceleration and are derived using calculus . I assume you haven't done any calculus yet?

OpenStudy (anonymous):

This is for a calculus paper

OpenStudy (cwrw238):

i see - then you should use calculus unfortunately i have to go now so someone else may help you

OpenStudy (anonymous):

I know but ai don't know how - that's why I need help with it :/

OpenStudy (anonymous):

Evaluate the velocity of a body and its distance moved, both after 2 minutes from it starts from rest, when its acceleration is given by a = 15 m/s^2